# Real-time 3D-aware Portrait Editing from a Single Image

### v1

@@ Abstract.
This work presents 3DPE, a practical tool that can efficiently edit a face image following given prompts, like reference images or text descriptions, in the 3D-aware manner. To this end, a lightweight module is distilled from a 3D portrait generator and a text-to-image model, which provide prior knowledge of face geometry and open-vocabulary editing capability, respectively. Such a design brings two compelling advantages over existing approaches. First, our system achieves real-time editing with a feedforward network (i.e., $\sim$0.04s per image), over 100$\times$ faster than the second competitor. Second, thanks to the powerful priors, our module could focus on the learning of editing-related variations, such that it manages to handle various types of editing simultaneously in the training phase and further supports fast adaptation to user-specified novel types of editing during inference (e.g., with $\sim$5min fine-tuning per case). The code, the model, and the interface will be made publicly available to facilitate future research.
3D-aware portrait, efficient editing
Figure 1. Photorealistic editing results produced by our proposed 3DPE, which allows users to perform 3D-aware portrait editing using image or text prompts. In comparison with baseline methods, such as InstructPix2Pix (Brooks et al., 2023)+Live3D (Trevithick et al., 2023) and PTI (Roich et al., 2022)+CLIP (Radford et al., 2021) (details are illustrated in Sec. 4), our approach accurately follows the guidance from reference prompts and maintains sufficiently better efficiency.
1. Introduction
@@
Inferring the geometry and appearance from a single-view portrait image has become mature and practical (Lin et al., 2022; Ko et al., 2023; Xie et al., 2023; Sun et al., 2022a; Wang et al., 2022; Trevithick et al., 2023; Gao et al., 2020), largely attributed to the utilization of priors in various 2D/3D generative models. However, only performing geometry reconstruction is insufficient. The significance of 3D editing of portraits, driven by user intentions, and the need for streamlined efficiency in the editing process has been steadily increasing. This is particularly crucial in real-world applications such as AR/VR, 3D telepresence, and video conferencing, where real-time editing is often essential. Consequently, a key question arises: How can we effectively address the challenge of attaining high-fidelity portrait editing while ensuring real-time efficiency?
@@
Traditional tools for 3D portrait editing (Deng et al., 2020) typically rely on template facial models (Blanz and Vetter, 2023; Paysan et al., 2009), which has limitations in handling the substantial geometry changes because it often overlooks precise features like hair and beard. Recent 3D GANs (Schwarz et al., 2020; Chan et al., 2021, 2022; Shi et al., 2022a; Xia and Xue, 2023) show remarkable capabilities in generating high-fidelity 3D portraits. They can serve as powerful generative priors for 3D portrait editing when coupled with GAN inversion techniques (Zhu et al., 2016, 2020; Shen et al., 2020b; Abdal et al., 2019, 2020). However, these methods may encounter issues related to geometry distortion or exhibit slow speed, and the editing is constrained by the limited latent attributes. Besides, 2D diffusion models can offer a strong prior with SDS loss (Poole et al., 2022) for editing purposes (Hertz et al., 2023). Nonetheless, they often require step-by-step optimization and become a major bottleneck for real-time applications.
@@
To this end, we present 3DPE, a real-time 3D-aware Portrait Editing method driven by user-defined prompts. As shown in Fig. 1, when provided with a single-view portrait image, our approach empowers a versatile range of editing through flexible instructions, including images and text. We leverage the powerful 3D prior from a 3D-aware face generator and achieve a high-fidelity 3D reconstruction of the portrait image. Subsequently, we distill editing knowledge from an open-vocabulary text-to-image model into a lightweight module integrated with the 3D-aware generator. This module, characterized by its minimal computational cost, allows our system to maintain real-time inference and excel in handling various types of editing, all the while ensuring good 3D consistency. An additional advantage is that our model supports customization through user-specified prompts with fast adaptation speed. This empowers users to build their own editing model at a minimal cost, enabling our system to cater to a broader audience.
@@
Our system achieves real-time 3D-aware portrait editing through the utilization of a feedforward network, with a processing time of 40ms on a standard customer GPU. Additionally, we present a comprehensive evaluation of our method using various prompts both quantitatively and qualitatively. Our design choices are also validated through comparisons with ablated variants of our method. Compared with baseline methods, our approach demonstrates superior 3D consistency, precise texture alignment, and a substantial improvement in inference time, as reflected by the evaluation metrics. We demonstrate the versatility of our method by showcasing its capacity to perform a wide variety of edits, including text and image prompts, on portrait images. The code, the model and the interface will be made publicly available.
@@
  • Our model can accommodate various control signals, including text and image prompts.
2. Related Work
2.1. Generative Face Priors
@@
Generative models aim at modeling the underlying distribution of the data, containing a wealth of prior knowledge. Recently, 3D GANs (Schwarz et al., 2020; Chan et al., 2021, 2022; Xu et al., 2022; Skorokhodov et al., 2022; Schwarz et al., 2022; Shi et al., 2022b, 2023; Pan et al., 2021; Gu et al., 2022; Or-El et al., 2022; Wang et al., 2023b) are mostly adopted to learn 3D faces from single-view image dataset. The interior rich domain-specific geometry priors enable various face-related applications such as image editing (Trevithick et al., 2023; Cai et al., 2022; Sun et al., 2022b, a). Large-scale diffusion models (Rombach et al., 2022), in contrast, encode knowledge of huge datasets and thus can provide general prior information. Such priors are broadly leveraged for tasks such as image editing (Meng et al., 2021; Zhang et al., 2023; Graikos et al., 2022; Brooks et al., 2023; Haque et al., 2023; Cao et al., 2023), customization (Ruiz et al., 2023; Hu et al., 2021; Liu et al., 2023a; Kumari et al., 2023), and video editing (Ouyang et al., 2023; Liu et al., 2023b; Ceylan et al., 2023; Qi et al., 2023; Wang et al., 2023a; Yang et al., 2023; Geyer et al., 2023; Chai et al., 2023). In our approach, we capitalize on the advantages of both the geometry prior derived from 3D GANs and the broader editing prior offered by large-scale text-to-image diffusion models, instead of relying solely on a single type of prior.
Figure 2. Overview of our method. We distill the prior in the diffusion model and 3D GAN for real-time 3D-aware editing. Our approach is fine-tuned from Live3D (Trevithick et al., 2023), where we extract features from the input portrait $\mathrm{I}$ using $\mathrm{E}_{high}(\cdot)$ and $\mathrm{E}_{low}(\cdot)$. The prompt embedding is generated with $\mathrm{E}_{p}(\cdot)$ and injected with the input features from $\mathrm{E}_{high}(\cdot)$ through a cross-attention mechanism. Our model is trained to mimic the output from the diffusion prior to acquire editing knowledge and enforce geometry constraints through triplane, multi-view images, and depth supervision from the 3D prior. In this context, InstructPix2Pix (Brooks et al., 2023) and Live3D serve as the diffusion and 3D prior, respectively. It’s noteworthy that only $\mathrm{E}_{p}(\cdot)$ and $\mathrm{E}_{t}(\cdot)$ are learnable during training, while all other parameters remain frozen.
2.2. Portrait Editing from a Single Image
@@
Although many methods work well for reconstructing (Guo et al., 2022; Gao et al., 2020) or generating (Chan et al., 2022) faces, editing has become a necessary interface to connect these methods with real-world applications. Previously, portrait editing given single-view image was mostly completed in 2D image space facilitated with GAN inversion (Shen et al., 2020b; Zhu et al., 2020; Abdal et al., 2019, 2020; Roich et al., 2022; Shen et al., 2020a; Xu et al., 2021; Bai et al., 2022; Härkönen et al., 2020), which enables fast editing by exploring the trajectories in the GAN’s latent space. Recent diffusion-based portrait editing (Brooks et al., 2023; Zhang et al., 2023) can support various editing types with texts as guidance. However, most of them are done in the 2D space, and thus there is no guarantee for the underlying 3D consistency. Therefore, 3D-aware portrait editing is crucial to achieve the goal. Some methods (Sun et al., 2022b; Lin et al., 2022; Sun et al., 2022a; Wang et al., 2022; Xie et al., 2023; Yin et al., 2023; Chang et al., 2023; Li et al., 2023a; Jiang et al., 2023) rely on the latent space of 3D GANs to perform editing through walking in the latent space, but are limited by the number of editable latent attributes. A few methods, such as Instruct-NeRF2NeRF (Haque et al., 2023), ClipFace (Aneja et al., 2023), and LENeRF (Hyung et al., 2023), attempt to leverage the large-scale open-vocabulary models (Radford et al., 2021; Rombach et al., 2022) to edit on 3D representations with texts as guidance but require heavy iterative refinement of the edited results. Similarly, InstructPix2NeRF (Li et al., 2023b) leverages text-to-image models and text prompts to edit the latent space of the GAN-based 3D generator. However, it is held back by the multi-step diffusion inference and the heavy structure of a GAN-based generator. Ours, in contrast, benefits from the lightweight design of the module and achieves real-time editing with ease. With such a design, our model can also adapt to novel editing prompts faster than previous methods. Moreover, texts cannot always illustrate the desired effects precisely, and sometimes an image prompt serves as a better guidance for image editing. Our method can support not only text prompts but also image prompts for editing, providing a more flexible and user-friendly interface.
3. Method
@@
Our approach takes a portrait image $\mathbf{I}$ and its camera pose c, obtained via the face pose estimator (Deng et al., 2019b). Additionally, the referenced prompts $\mathbf{P}$ (images or texts) serve as the editing instructions. The outcome of our model is an edited version of the 3D portraits characterized by NeRF in accordance with the prompts. The edited image is denoted as $\mathbf{I}_{p}$. Within our framework, we achieve this objective by leveraging the knowledge of 3D GANs and an open-vocabulary text-to-image model, which is distilled into a lightweight module. Following distillation, our method enables real-time 3D portrait editing and efficient adaptation to user-specified prompts.
@@
In Sec. 3.1, we offer an illustration of 3D GANs and diffusion models. The process of distilling these priors into a lightweight module is outlined in Sec. 3.2. Finally, the details for model training and inference, along with the fast adaptation for novel prompts, are presented in Sec. 3.3.
3.1. Preliminary
@@
3D GAN Prior for Portrait Reconstruction. The 3D-aware GANs showcase the ability to synthesize photorealistic 3D images using a collection of single-view images. Notably, EG3D (Chan et al., 2022) introduces an efficient triplane 3D representation, demonstrating high-quality 3D-aware image rendering. Once trained, the generator of EG3D can be applied for single-image 3D reconstruction via GAN inversion. However, existing 3D GAN inversion methods often encounter geometry distortion or exhibit slow inference speed. To tackle these challenges, we utilize Live3D (Trevithick et al., 2023), a state-of-the-art single-image portrait 3D reconstruction model built upon EG3D, preserving geometry quality while ensuring real-time performance. Live3D employs a two-branch encoder, $\mathrm{E}_{high}(\cdot)$ and $\mathrm{E}_{low}(\cdot)$, to extract different resolution features from the input image. It then utilizes a ViT-based $\mathrm{E}_{t}(\cdot)$ decoder (Dosovitskiy et al., 2020) to transform the fused encoder features into the triplane, $\mathbf{T}$:
(1) $\mathbf{T}=\mathrm{E}_{t}(\mathrm{E}_{high}(\mathbf{I}),\mathrm{E}_{low}(\mathbf{I}))$
@@
Diffusion Prior for Portrait Editing. Only performing 3D reconstruction is insufficient as our goal is to edit portrait images. Large-scale diffusion models (Rombach et al., 2022), trained on vast text-image pairs, can synthesize realistic photos with text input, offering a powerful editing prior. Despite their effectiveness for face images, the editing process is often slow due to step-by-step optimization or multiple inferences in diffusion models. Moreover, obtaining a large amount of high-quality paired data with multi-view consistency from the diffusion model is extremely challenging, resulting in the application of 3D-aware editing from a single image impractical. Therefore, our goal is to distill editing knowledge from the diffusion model, integrate it with the 3D prior in Live3D into a lightweight module, and employ it for real-time portrait editing.
3.2. Distilling Priors into a Lightweight Module
Our system aims to perform real-time 3D-aware editing for various prompts for the single-view portrait. Thus, it needs powerful 3D knowledge for geometry reconstruction and editing prior to handling various control signals. We leverage the strengths of both 3D GANs and diffusion models. In the following, we provide detailed presentations on how the knowledge of these two types of models is distilled into a lightweight module.
Feature Representation in Live3D. We carefully study the Live3D model and discover that, as a two-branch triplane-based 3D reconstruction model, the low-resolution and high-resolution features from the Live3D encoder $\mathrm{E}_{low}(\cdot)$ and $\mathrm{E}_{high}(\cdot)$ tend to learn various levels of information without explicit guidance. We conduct a study by disabling one of the branches and inferring the reconstructed images. As illustrated in Fig. 3, when the high-resolution encoder $\mathrm{E}_{high}(\cdot)$ is disabled, the inference image retains a similar structure but loses its detailed appearance. Conversely, when the low-resolution encoder $\mathrm{E}_{low}(\cdot)$ is disabled, the reconstructed portraits preserve some of the texture from the input but struggle to capture the geometry. Based on this analysis, the features of the two branches separately model low-frequency and high-frequency information. This insight motivates our model design in terms of how to distill knowledge from the diffusion model and 3D GANs.

### v2

@@ Abstract
This work presents 3DPE, a practical method that can efficiently edit a face image following given prompts, like reference images or text descriptions, in a 3D-aware manner. To this end, a lightweight module is distilled from a 3D portrait generator and a text-to-image model, which provide prior knowledge of face geometry and superior editing capability, respectively. Such a design brings two compelling advantages over existing approaches. First, our system achieves real-time editing with a feedforward network (i.e., $\sim$0.04s per image), over 100$\times$ faster than the second competitor. Second, thanks to the powerful priors, our module could focus on the learning of editing-related variations, such that it manages to handle various types of editing simultaneously in the training phase and further supports fast adaptation to user-specified customized types of editing during inference (e.g., with $\sim$5min fine-tuning per style). The code, the model, and the interface will be made publicly available to facilitate future research.
Keywords:
3D-aware portrait Efficient editing
1 Introduction
@@
Inferring the geometry and appearance from a single-view portrait image has become mature and practical [42, 37, 74, 65, 72, 68, 22], largely attributed to the utilization of priors in various 2D/3D generative models. However, only performing geometry reconstruction is insufficient. The significance of 3D editing of portraits, driven by user intentions, and the need for streamlined efficiency in the editing process has been steadily increasing. This is particularly crucial in real-world applications such as AR/VR, 3D telepresence, and video conferencing, where real-time editing is often essential. Consequently, a key question arises: How can we effectively address the challenge of attaining high-fidelity portrait editing while ensuring real-time efficiency?
Figure 1: Photorealistic editing results produced by our proposed 3DPE, which allows users to perform 3D-aware portrait editing using image or text prompts. In comparison with baseline methods, such as InstructPix2Pix (IP2P) [7]+Live3D [68] and PTI [54]+StyleCLIP [49] (details are illustrated in Sec. 4), our approach accurately follows the guidance from reference prompts and maintains sufficiently better efficiency.
@@
Traditional methods for 3D portrait editing [17] typically rely on template facial models [6, 50], which has limitations in handling the substantial geometry changes because it often overlooks precise features like hair and beard. Recent 3D GANs [57, 14, 13, 61, 73] show remarkable capabilities in generating high-fidelity 3D portraits. They can serve as powerful generative priors for 3D portrait editing when coupled with GAN inversion techniques [85, 84, 60, 2, 3]. However, these methods may encounter issues related to geometry distortion or exhibit slow speed, and the editing is constrained by the limited latent attributes. Besides, 2D diffusion models can offer a strong prior with Score Distillation Sampling (SDS) loss [51] for editing purposes [30]. Nonetheless, they often require step-by-step optimization and become a major bottleneck for real-time applications.
@@
To this end, we present 3DPE, a real-time 3D-aware Portrait Editing method driven by user-defined prompts. As shown in Fig. 1, when provided with a single-view portrait image, our approach empowers a versatile range of editing styles with one model through flexible instructions, including images and texts. We leverage the powerful 3D prior from a 3D-aware face generator and achieve a high-fidelity 3D reconstruction of the portrait image. Subsequently, we distill editing knowledge from a text-guided image editing model into a lightweight module integrated with the 3D-aware generator. This module, characterized by its minimal computational cost, allows our system to maintain real-time inference and excel in handling various types of editing, all the while ensuring good 3D consistency. An additional advantage is that our model supports customization through user-specified prompts with fast adaptation speed. This empowers users to build their own editing model at a minimal cost, enabling our system to cater to a broader audience.
@@
Our system achieves real-time 3D-aware portrait editing through the utilization of a feedforward network, with a processing time of 40ms on a standard consumer GPU. Additionally, we present a comprehensive evaluation of our method using various prompts both quantitatively and qualitatively. Our design choices are also validated through comparisons with ablated variants of our method. Compared with baseline methods, our approach demonstrates superior 3D consistency, precise texture alignment, and a substantial improvement in inference time, as reflected by the evaluation metrics. We demonstrate the versatility of our method by showcasing its capacity to perform a wide variety of edits, including text and image prompts, on portrait images. The code, the model, and the interface will be made publicly available.
@@
  • Our framework can accommodate various control signals, including text and image prompts.
2 Related Work
2.1 Generative Face Priors
@@
Generative models aim at modeling the underlying distribution of the data, containing a wealth of prior knowledge. Recently, 3D GANs [57, 14, 13, 75, 64, 58, 63, 62, 48, 25, 46, 69] are mostly adopted to learn 3D faces from single-view image dataset. The interior rich domain-specific geometry priors enable various face-related applications such as image editing [68, 9, 66, 65, 83, 77] and domain adaptation [34, 36, 80, 1]. Large-scale diffusion models [55], in contrast, encode knowledge of huge datasets and thus can provide general prior information. Such priors are broadly leveraged for tasks such as image editing [45, 81, 24, 7, 10], customization [56, 31, 44, 38], and video editing [47, 43, 11, 52, 71, 78, 23, 12]. In our approach, we capitalize on the advantages of both the geometry prior derived from 3D GANs and the broader editing prior offered by large-scale text-to-image diffusion models, instead of relying solely on a single type of prior.
2.2 Portrait Editing from a Single Image
@@
Although many methods work well for reconstructing [26, 22] or generating [13] faces, editing has become a necessary interface to connect these methods with real-world applications. Previously, portrait editing given a single-view image was mostly completed in 2D image space facilitated with GAN inversion [60, 84, 2, 3, 54, 59, 76, 5, 28], which enables fast editing by exploring the trajectories in the GAN’s latent space. Recent diffusion-based portrait editing [7, 81] can support various editing types with texts as guidance. However, most of them are done in the 2D space, and thus there is no guarantee for the underlying 3D consistency. Therefore, 3D-aware portrait editing is crucial to achieve the goal. Some methods [66, 42, 65, 72, 74, 79, 15, 39, 33] rely on the latent space of 3D GANs to perform editing through walking in the latent space, but are limited by the number of editable latent attributes. A few methods, such as Instruct-NeRF2NeRF [27], ClipFace [4], and LENeRF [32], attempt to leverage the large-scale models [53, 55] to edit on 3D representations with texts as guidance but require heavy iterative refinement of the edited results. Similarly, InstructPix2NeRF [40] leverages text-to-image models and text prompts to edit within the latent space of the GAN-based 3D generator. However, it is held back by the multi-step diffusion inference and the heavy structure of a GAN-based generator. Ours, in contrast, benefits from the lightweight design of the module and achieves real-time editing with ease. With such a design, our model can also adapt fast to the user-defined editing prompts. Moreover, texts cannot always illustrate the desired effects precisely, and sometimes an image prompt serves as better guidance for image editing [20, 56]. Our method can support not only text prompts but also image prompts for editing, providing a more flexible and user-friendly interface.
3 Method
@@
Our approach takes a portrait image $\mathbf{I}$ and its camera pose c, obtained via the face pose estimator [18]. Additionally, the referenced prompts $\mathbf{P}$ (images or texts) serve as the editing instructions. The outcome of our model is an edited version of the 3D portraits characterized by NeRF in accordance with the prompts. The edited image is denoted as $\mathbf{I}_{p}$. Within our framework, we achieve this objective by leveraging the knowledge of 3D GANs and a text-guided image editing model, which is distilled into a lightweight module. Following distillation, our method enables real-time 3D portrait editing and efficient adaptation to user-specified prompts.
@@
In Sec. 3.1, we offer an illustration of 3D GANs and diffusion models. The process of distilling these priors into a lightweight module is outlined in Sec. 3.2. Finally, the details for model training and inference, along with the fast adaptation for customized prompts, are presented in Sec. 3.3.
Figure 2: Overview of our method. We distill the prior in the diffusion model and 3D GAN for real-time 3D-aware editing. Our approach is fine-tuned from Live3D [68], where we extract features from the input portrait $\mathrm{I}$ using $\mathrm{E}_{high}(\cdot)$ and $\mathrm{E}_{low}(\cdot)$. The prompt embedding is generated with $\mathrm{E}_{p}(\cdot)$ and injected with the input features from $\mathrm{E}_{high}(\cdot)$ through a cross-attention mechanism. Our model is trained to mimic the output from the diffusion prior to acquire editing knowledge and enforce geometry constraints through triplane, multi-view images, and depth supervision from the 3D prior. In this context, InstructPix2Pix [7] and Live3D serve as the diffusion and 3D prior, respectively. It’s noteworthy that only $\mathrm{E}_{p}(\cdot)$ and $\mathrm{E}_{t}(\cdot)$ are learnable during training, while all other parameters remain frozen.
3.1 Preliminary
@@
3D GAN Prior for Portrait Reconstruction. The 3D-aware GANs showcase the ability to synthesize photorealistic 3D images using a collection of single-view images. Notably, EG3D [13] introduces an efficient triplane 3D representation, demonstrating high-quality 3D-aware image rendering. Once trained, the generator of EG3D can be applied for single-image 3D reconstruction via GAN inversion. However, existing 3D GAN inversion methods often encounter geometry distortion or exhibit slow inference speed. To tackle these challenges, we utilize Live3D [68], a state-of-the-art single-image portrait 3D reconstruction model built upon EG3D, preserving geometry quality while ensuring real-time performance. Live3D employs a two-branch encoder, $\mathrm{E}_{high}(\cdot)$ and $\mathrm{E}_{low}(\cdot)$, to extract different resolution features from the input image. It then utilizes a ViT-based $\mathrm{E}_{t}(\cdot)$ decoder [19] to transform the fused encoder features into the triplane $\mathbf{T}$:
$\mathbf{T}=\mathrm{E}_{t}(\mathrm{E}_{high}(\mathbf{I}),\mathrm{E}_{low}(\mathbf{I})).$ (1)
@@
Diffusion Prior for Portrait Editing. Only performing 3D reconstruction is insufficient as our goal is to edit portrait images. Large-scale diffusion models [55], trained on vast text-image pairs, can synthesize realistic photos with text input, offering powerful editing prior. Despite their effectiveness for face images, the editing process is often slow due to step-by-step optimization or multiple inferences in diffusion models. Moreover, obtaining a large amount of high-quality paired data with multi-view consistency from the diffusion model is extremely challenging, making the application of 3D-aware editing from a single image impractical. Therefore, our goal is to distill editing knowledge from the diffusion model, integrate it with the 3D prior in Live3D into a lightweight module, and employ it for real-time portrait editing.
Figure 3: Disentanglement in Live3D features. We separately disable the features from $\mathrm{E}_{high}(\cdot)$ and $\mathrm{E}_{low}(\cdot)$ to infer the reconstructed image. Without $\mathrm{E}_{high}(\cdot)$, the output retains the coarse structure but loses its appearance. Conversely, when $\mathrm{E}_{low}(\cdot)$ is deactivated, the reconstructed portraits preserve the texture (such as the blue and purple reflection on the glasses) but fail to capture the geometry.
3.2 Distilling Priors into a Lightweight Module
Our method aims to perform real-time 3D-aware editing for various prompts for the single-view portrait. Thus, it needs powerful 3D knowledge for geometry reconstruction and editing prior to handling various control signals. We leverage the strengths of both 3D GANs and diffusion models. In the following, we provide detailed presentations on how the knowledge of these two types of models is distilled into a lightweight module.
Feature Representation in Live3D. We carefully study the Live3D model and discover that, as a two-branch triplane-based 3D reconstruction model, the low-resolution and high-resolution features from the Live3D encoder $\mathrm{E}_{low}(\cdot)$ and $\mathrm{E}_{high}(\cdot)$ tend to learn various levels of information without explicit guidance. We conduct a study by disabling one of the branches and inferring the reconstructed images. As illustrated in Fig. 3, when the high-resolution encoder $\mathrm{E}_{high}(\cdot)$ is disabled with random weights, the inference image retains a similar structure but loses its detailed appearance. Conversely, when the low-resolution encoder $\mathrm{E}_{low}(\cdot)$ is disabled, the reconstructed portraits preserve some of the texture from the input but struggle to capture the geometry. Based on this analysis, the features of the two branches separately model low-frequency and high-frequency information. This insight motivates our model design in terms of how to distill knowledge from the diffusion model and 3D GANs.